# Implicit Bias of Mirror Descent for Shallow Neural Networks in Univariate Regression

### v1

@@ A.3 Function description of the implicit bias
Next we aim to characterize the solution to (41) in the function space:
@@
Consider a two layer absolute value network (1) with $n$ hidden units, where we assume $n$ is sufficiently large. Consider parameter initialization (2) and, specifically, let $p_{\mathcal{B}}(b)$ denote the density function for random variable $\mathcal{B}$. Consider any finite training data set $\{(x_{i},y_{i})\}_{i=1}^{m}$ that satisfies $x_{i}\neq x_{j}$ when $i\neq j$ and $\{x_{i}\}_{i=1}^{m}\subset\operatorname{supp}(p_{\mathcal{B}})$, and consider mirror flow (4) with a potential function $\Phi$ satisfying Assumption 1 and learning rate $\eta=\Theta(1/n)$. Then, with high probability over the random parameter initialization, there exist constants $C_{1},C_{2}>0$ such that for any $t\geq 0$,
$\|\theta(t)-\hat{\theta}\|_{\infty}\leq C_{1}n^{-1},\ \lim_{n\rightarrow\infty}\|H(t)-H(0)\|_{2}=0,\ \|\mathbf{f}(t)-\mathbf{y}\|_{2}\leq e^{-\eta_{0}C_{2}t}\|\mathbf{f}(0)-\mathbf{y}\|_{2}.$
Moreover, letting $\theta(\infty)=\lim_{t\rightarrow\infty}\theta(t)$, we have for any given $x\in\mathbb{R}$, that $\lim_{n\rightarrow\infty}|f(x,\theta(\infty))-f(x,\hat{\theta})-\bar{h}(x)|=0$, where $\bar{h}(\cdot)$ is the solution to the following variational problem:
@@ B.1 Linearization
$H(t)=\frac{1}{n}J_{\theta}\mathbf{f}(t)(\nabla^{2}\Phi(\theta))^{-1}J_{\theta}\mathbf{f}(t)^{T},$
@@
Consider a single-hidden-layer univariate network (1) with ReLU or absolute value activation function and parameter initialization (2). Assume $\mathcal{B}$ has a compact support. Assume the number of hidden neurons is sufficiently large. Given any finite training data set $\{(x_{i},y_{i})\}_{i=1}^{m}$ that satisfies $x_{i}\neq x_{j}$ when $i\neq j$ and $\{x_{i}\}_{i=1}^{m}\subset\operatorname{supp}(p_{\mathcal{B}})$, consider mirror flow (4) with learning rate $\eta=\eta_{0}/n$, $\eta_{0}>0$ and potential function $\Phi$. Assume $\Phi$ can be written in the form of $\Phi(\theta)=\frac{1}{n^{2}}\sum_{k=1}^{3n+1}\phi\big{(}n(\theta_{k}-\hat{\theta}_{k})\big{)}$, where $\phi$ is a real-valued $C^{3}$-smooth function on $\mathbb{R}$ and has strictly positive second derivatives everywhere. Assume there exists constant $C_{1}>0$ satisfying
$\frac{\kappa_{1}(C_{1})}{\kappa_{2}(C_{1})}\geq\frac{8\sqrt{m}(B_{x}+1)K\|\mathbf{f}(0)-\mathbf{y}\|_{2}}{\lambda_{0}}\cdot\frac{1}{C_{1}},$ (69)
where $\lambda_{0}=\operatorname{plim}_{n\rightarrow\infty}\lambda_{\min}(\frac{1}{n}J_{\theta}\mathbf{f}(0)J_{\theta}\mathbf{f}(0)^{T})$, $m$ is the size of the training data set, $B_{x}=\max_{i\in[m]}|x_{i}|$, and $K$ is a constant determined by $\mathcal{B}$ and $\mathcal{A}$. Then with high probability over random initialization,
@@
$\|\mathbf{f}(t)-\mathbf{y}\|_{2}\leq\exp\Big{(}-\frac{\kappa_{1}(C_{1})\lambda_{0}}{4}\eta_{0}t\Big{)}\|\mathbf{f}(0)-\mathbf{y}\|_{2},\ \forall t\geq 0.$
@@
[Proof of Proposition 24] To ease the notation, let $J(t)=J_{\theta}\mathbf{f}(t)$ and $Q(t)=(\nabla^{2}\Phi(\theta(t)))^{-1}$. For $i\in[m]$ and $k\in[n]$, let $\sigma_{ik}(t)=\sigma(w_{k}(t)x_{i}-b_{k}(t))$ and $\sigma^{\prime}_{ik}(t)=\sigma^{\prime}(w_{k}(t)x_{i}-b_{k}(t))$.
@@
$\begin{dcases}I_{1}=\frac{1}{n}\sum_{k=1}^{n}|\sigma_{ik}(t)\sigma_{jk}(t)-\sigma_{ik}(0)\sigma_{jk}(0)|;\\ I_{2}=|x_{i}x_{j}|\frac{1}{n}\sum_{k=1}^{n}|a_{k}(t)^{2}\sigma^{\prime}_{ik}(t)\sigma^{\prime}_{jk}(t)-a_{k}(0)^{2}\sigma^{\prime}_{ik}(0)\sigma^{\prime}_{jk}(0)|;\\ I_{3}=\frac{1}{n}\sum_{k=1}^{n}\Big{|}a_{k}(t)^{2}\sigma^{\prime}_{ik}(t)\sigma^{\prime}_{jk}(t)-a_{k}(0)^{2}\sigma^{\prime}_{ik}(0)\sigma^{\prime}_{jk}(0)|.\end{dcases}$ (71)
@@
$\frac{\mathrm{d}}{\mathrm{dt}}\|\mathbf{f}(t)-\mathbf{y}\|_{2}^{2}\leq-2\eta_{0}\frac{\kappa_{1}(C_{1})\lambda_{0}}{4}\|\mathbf{f}(t)-\mathbf{y}\|_{2}^{2}$
@@
$\|\mathbf{f}(t)-\mathbf{y}\|_{2}\leq\exp(-\frac{\kappa_{1}(C_{1})\lambda_{0}}{4}\eta_{0}t)\|\mathbf{f}(0)-\mathbf{y}\|_{2}.$ (74)
@@
$\|J_{a_{k}}\mathbf{f}(t)\|_{2}\leq\sqrt{m}\|J_{a_{k}}\mathbf{f}\|_{\infty}\leq\sqrt{m}(B_{x}|w_{k}(t)|+|b_{k}(t)|)\leq 2\sqrt{m}K(B_{x}+1).$ (75)
@@
$|\frac{\mathrm{d}}{\mathrm{dt}}a_{k}(t)|$ $\leq\frac{\eta_{0}}{n}\|Q(t)\|_{2}\|J_{a_{k}}\mathbf{f}(t)\|_{2}\|\mathbf{f}(t)-\mathbf{y}\|_{2}$ (76)
$\leq\frac{\eta_{0}}{n}\cdot\kappa_{2}(C_{1})\cdot 2\sqrt{m}K(B_{x}+1)\cdot\exp(-\frac{\kappa_{1}(C_{1})\lambda_{0}}{4}\eta_{0}t)\|\mathbf{f}(0)-\mathbf{y}\|_{2}.$
@@
$<\frac{8\kappa_{2}(C_{1})\sqrt{m}K(B_{x}+1)\|\mathbf{f}(0)-\mathbf{y}\|_{2}}{\kappa_{1}(C_{1})\lambda_{0}}n^{-1}$
@@
$\begin{dcases}\|J_{w_{k}}\mathbf{f}(t)\|_{2}\leq\sqrt{m}B_{x}|a_{k}(t)|\leq 2\sqrt{m}B_{x}K,\quad\forall k\in[n];\\ \|J_{b_{k}}\mathbf{f}(t)\|_{2}\leq\sqrt{m}|a_{k}(t)|\leq 2\sqrt{m}K,\quad\forall k\in[n];\\ \|J_{d}\mathbf{f}(t)\|_{2}=\sqrt{m}.\end{dcases}$ (78)
@@
$\|\tilde{\theta}(t)-\hat{\theta}\|_{\infty}\leq C_{1}n^{-1},\quad\|\tilde{\mathbf{f}}(t)-\mathbf{y}\|_{2}\leq\exp(-\frac{\kappa_{1}(C_{1})\lambda_{0}}{4}\eta_{0}t)\|\mathbf{f}(0)-\mathbf{y}\|_{2}.$
@@
$\kappa_{2}(C_{1})\kappa_{3}(C_{1})\Big{(}\lambda_{0}\frac{\kappa_{2}(C_{1})}{\kappa_{1}(C_{1})}+10m(B_{x}+1)^{2}K^{2}\Big{(}\frac{\kappa_{2}(C_{1})}{\kappa_{1}(C_{1})}\Big{)}^{2}\Big{)}\leq\frac{(\lambda_{0})^{2}}{16(B_{x}+1)K\sqrt{m}\|\mathbf{f}(0)-\mathbf{y}\|_{2}}.$ (79)
Then for any $x\in\mathbb{R}$ and with high probability over initialization,
@@
For any fixed $C>0$, consider $t_{1}=\inf\{t\geq 0\colon\iota(t)>C\}$. Now we aim to prove $t_{1}=+\infty$. For the sake of contradiction, assume $t_{1}<+\infty$. Then we have for $t\in[0,t_{1}]$, $\iota(t)\leq C$ and especially, $\iota(t_{1})=C$.
@@
Next we bound the difference between $H(t)$ and $\tilde{H}(t)$. To simplify the notation, we let $\sigma_{ik}(t)=\sigma(w_{k}(t)x_{i}-b_{k}(t))$ for $i\in[m]$ and $k\in[n]$. For $i,j\in[m]$ we have the following decomposition:
$|H_{ij}(t)-\tilde{H}_{ij}(0)|\leq I_{1}+I_{2}+I_{3},$ (80)
@@
$\begin{dcases}I_{1}=\frac{1}{n}\sum_{k=1}^{n}\Big{|}\frac{\sigma_{ik}(t)\sigma_{jk}(t)}{\phi^{\prime\prime}(n(a_{k}(t)-a_{k}(0)))}-\frac{\sigma_{ik}(0)\sigma_{jk}(0)}{\phi^{\prime\prime}(n(\tilde{a}_{k}(t)-a_{k}(0)))}\Big{|};\\ I_{2}=\frac{1}{n}\sum_{k=1}^{n}\Big{(}|x_{i}x_{j}|\Big{|}\frac{a_{k}(t)^{2}\sigma^{\prime}_{ik}(t)\sigma^{\prime}_{jk}(t)}{\phi^{\prime\prime}(n(w_{k}(t)-w_{k}(0)))}\Big{|}+\Big{|}\frac{a_{k}(t)^{2}\sigma^{\prime}_{ik}(t)\sigma^{\prime}_{jk}(t)}{\phi^{\prime\prime}(n(b_{k}(t)-b_{k}(0)))}\Big{|}\Big{)}\\ I_{3}=\frac{1}{n}\Big{|}\frac{1}{\phi^{\prime\prime}(n(d(t)-d(0)))}\Big{|}.\end{dcases}$ (81)

### v2

@@ A.3 Function description of the implicit bias
Assume the input dimension is one, i.e., $d=1$. Next we aim to characterize the solution to (42) in the function space:
@@
Consider a two layer absolute value network (1) with $d$ input units and $n$ hidden units, where we assume $n$ is sufficiently large. Consider parameter initialization (2) and, specifically, let $p_{\mathcal{B}}(b)$ denote the density function for random variable $\mathcal{B}$. Consider any finite training data set $\{(x_{i},y_{i})\}_{i=1}^{m}$ that satisfies $x_{i}\neq x_{j}$ when $i\neq j$ and $\{\|x_{i}\|_{2}\}_{i=1}^{m}\subset\operatorname{supp}(p_{\mathcal{B}})$, and consider mirror flow (4) with a potential function $\Phi$ satisfying Assumption 1 and learning rate $\eta=\Theta(1/n)$. Then, there exist constants $C_{1},C_{2}>0$ such that with high probability over the random parameter initialization, for any $t\geq 0$,
$\|\theta(t)-\hat{\theta}\|_{\infty}\leq C_{1}n^{-1},\ \lim_{n\rightarrow\infty}\|H(t)-H(0)\|_{2}=0,\ \|\boldsymbol{f}(t)-\boldsymbol{y}\|_{2}\leq e^{-\eta_{0}C_{2}t}\|\boldsymbol{f}(0)-\boldsymbol{y}\|_{2}.$
Moreover, letting $\theta(\infty)=\lim_{t\rightarrow\infty}\theta(t)$, we have for any given $x\in\mathbb{R}^{d}$, that $\lim_{n\rightarrow\infty}|f(x,\theta(\infty))-f(x,\theta_{\mathrm{GF}}(\infty))|=0$, where $\theta_{\mathrm{GF}}(\infty)$ denotes the limiting point of gradient flow, i.e., mirror flow (4) with $\Phi(\theta)=\|\theta\|_{2}^{2}$, on the same training data and initial parameter.
Assuming univariate input data, i.e., $d=1$, we have for any given $x\in\mathbb{R}$, that $\lim_{n\rightarrow\infty}|f(x,\theta(\infty))-f(x,\hat{\theta})-\bar{h}(x)|=0$, where $\bar{h}(\cdot)$ is the solution to the following variational problem:
@@
A.4 Theorem 2 for uncentered potential
In this subsection, we characterize the implicit bias of mirror flow with “uncentered” and unscaled potential functions that satisfying the following assumptions.
Assumption 25.
The potential function $\Phi\colon\mathbb{R}^{3n+1}\rightarrow\mathbb{R}$ satisfies: (i) $\Phi$ can be written in the form $\Phi(\theta)=\sum_{k=1}^{3n+1}\phi(\theta_{k})$, where $\phi$ is a real-valued function on $\mathbb{R}$; (ii) $\Phi$ is twice continuously differentiable; (iii) $\nabla^{2}\Phi(\theta)$ is positive definite for all $\theta\in\mathbb{R}^{3n+1}$.
Theorem 26 (Theorem 2 for uncentered potentials).
Consider a two layer ReLU network (1) with $d$ input units and $n$ hidden units, where we assume $n$ is sufficiently large. Consider parameter initialization (2) and assume $\mathcal{B}$ is compactly supported on $[-B,B]$ for some $B>0$. Let $p_{\mathcal{B}}(b)$ denote the density function for random variable $\mathcal{B}$. Consider any finite training data set $\{(x_{i},y_{i})\}_{i=1}^{m}$ that satisfies $x_{i}\neq x_{j}$ when $i\neq j$ and $\{\|x_{i}\|_{2}\}_{i=1}^{m}\subset\operatorname{supp}(p_{\mathcal{B}})$, and consider mirror flow (4) with a potential function $\Phi$ satisfying Assumption 25 and learning rate $\eta=\Theta(1/n)$. Then, there exist constants $C_{1},C_{2}>0$ such that with high probability over the random parameter initialization, for any $t\geq 0$,
$\|\theta(t)-\hat{\theta}\|_{\infty}\leq C_{1}n^{-1/2},\ \lim_{n\rightarrow\infty}\|H(t)-H(0)\|_{2}=0,\ \|\boldsymbol{f}(t)-\boldsymbol{y}\|_{2}\leq e^{-\eta_{0}C_{2}t}\|\boldsymbol{f}(0)-\boldsymbol{y}\|_{2}.$ (70)
Moreover, letting $\theta(\infty)=\lim_{t\rightarrow\infty}\theta(t)$, we have for any given $x\in\mathbb{R}^{d}$, that $\lim_{n\rightarrow\infty}|f(x,\theta(\infty))-f(x,\theta_{\mathrm{GF}}(\infty))|=0$, where $\theta_{\mathrm{GF}}(\infty)$ denotes the limiting point of gradient flow, i.e., mirror flow (4) with $\Phi(\theta)=\|\theta\|_{2}^{2}$, on the same training data and initial parameter.
Assuming univariate input data, i.e., $d=1$, we have for any given $x\in\mathbb{R}$, that $\lim_{n\rightarrow\infty}|f(x,\theta(\infty))-f(x,\hat{\theta})-\bar{h}(x)|=0$, where $\bar{h}(\cdot)$ is the solution to the following variational problem:
where $\text{where }\begin{dcases}\mathcal{G}_{1}(h)=\int_{\operatorname{supp}(p_{\mathcal{B}})}\frac{\left(h^{\prime\prime}(x)\right)^{2}}{p_{\mathcal{B}}(x)}\mathrm{d}x\\ \mathcal{G}_{2}(h)=(h^{\prime}(+\infty)+h^{\prime}(-\infty))^{2}\\ \mathcal{G}_{3}(h)=\frac{1}{\mathbb{E}[\mathcal{B}^{2}]}(B(h^{\prime}(+\infty)-h^{\prime}(-\infty))-(h(B)+h(-B)))^{2}.\end{dcases}$
Here $\mathcal{F}_{\mathrm{ReLU}}=\{h(x)=\int\alpha(w,b)[wx-b]_{+}\mathrm{d}\mu\colon\alpha\in C(\operatorname{supp}(\mu)),\alpha\text{ is uniformly continuous}\}$, $h^{\prime}(+\infty)=\lim_{x\rightarrow+\infty}h^{\prime}(x)$, and $h^{\prime}(-\infty)=\lim_{x\rightarrow-\infty}h^{\prime}(x)$.
The proof of Theorem 26 mirrors the proof of Theorem 2, which we discussed in previous subsections. For brevity, below we only provide the proof sketch and highlight the differences brought by the change in assumptions.
First, we show Lemma 13 holds for uncentered potentials. Notice that
$H_{ij}(0)$ $=\frac{1}{n}\sum_{k=1}^{3n+1}\nabla_{\theta_{k}}f(x_{i};\hat{\theta})\frac{1}{\phi^{\prime\prime}(\hat{\theta}_{k})}\nabla_{\theta_{k}}f(x_{j};\hat{\theta})$
$=\frac{1}{n}\sum_{k=1}^{n}\frac{1}{\phi^{\prime\prime}(\hat{a}_{k})}\sigma(\hat{w}_{k}^{T}x_{i}-\hat{b}_{k})\sigma(\hat{w}_{k}^{T}x_{j}-\hat{b}_{k})+\frac{1}{n\phi^{\prime\prime}(\hat{d})}$
$\quad+\frac{1}{n}\sum_{k=1}^{n}\hat{a}_{k}^{2}\sigma^{\prime}(\hat{w}_{k}^{T}x_{i}-\hat{b}_{k})\sigma^{\prime}(\hat{w}_{k}^{T}x_{j}-\hat{b}_{k})\Big{(}\frac{1}{\phi^{\prime\prime}(\hat{b}_{k})}+\sum_{r=1}^{d}\frac{x_{i,r}x_{j,r}}{\phi^{\prime\prime}(\hat{w}_{k,r})}\Big{)}.$
By sub-Gaussian concentration inequality, we have $\max_{k\in[n]}|\hat{a}_{k}|=O_{p}(\sqrt{(\log n)/n})$. By noticing that $\max_{k\in[n],i\in[m]}\sigma(\hat{w}_{k}^{T}x_{i}-\hat{b}_{k})=O(1)$ since $\mathcal{W}$ and $\mathcal{B}$ are bounded, we have
[…]
Next, we show Proposition 14 and Proposition 16 hold. Note the key ingredients which are needed in proving these two results and which relate to the potentials are: (i) there exists a constant $\beta>0$ such that with high probability, $\|Q(0)\|_{2}\leq\beta$ and $\|Q(t)\|_{2}\leq\beta$ if $\|\theta(t)-\theta(0)\|_{\infty}=o_{p}(1)$; and (ii) $\|Q(t)-Q(0)\|_{2}=o_{p}(1)$ if $\|\theta(t)-\theta(0)\|_{\infty}=o_{p}(1)$, where we use the same notation as in Proposition 14. We verify these two properties for uncentered potentials. Since $\mathcal{B}$ is compactly supported, we have with high probability, $\max_{k\in[3n+1]}|\theta_{k}(0)|<C$ for some constant $C$ independent of $n$. Since $1/\phi^{\prime\prime}(\cdot)$ is uniformly continuous on $[-2C,2C]$, we have
$\|Q(t)-Q(0)\|_{2}=\max_{k\in[3n+1]}\Big{|}\frac{1}{\phi^{\prime\prime}(\theta_{k}(t))}-\frac{1}{\phi^{\prime\prime}(\theta_{k}(0))}\Big{|}=o_{p}(1).$
given $\|\theta(t)-\theta(0)\|_{\infty}=o_{p}(1)$. Meanwhile, notice that
$\|Q(t)\|_{2}=\max_{k\in[3n+1]}\frac{1}{\phi^{\prime\prime}(\theta_{k}(t))}\leq\max_{k\in[3n+1]}\frac{1}{\phi^{\prime\prime}(\theta_{k}(0))}+\Big{|}\frac{1}{\phi^{\prime\prime}(\theta_{k}(t))}-\frac{1}{\phi^{\prime\prime}(\theta_{k}(0))}\Big{|}.$
By the continuity of $\phi^{\prime\prime}$ we have $\|Q(t)\|_{2}$ can be bounded by some constant $\beta>0$ with high probability, given $\|\theta(t)-\theta(0)\|_{\infty}=o_{p}(1)$.
Then, we show Proposition 17 holds. Notice that for uncentered potentials,
$D_{\Phi}(\boldsymbol{a},\boldsymbol{a}(0))$ $=\sum_{k\in[n]}\Big{(}\phi(a_{k})-\phi(a_{k}(0))-\phi^{\prime}(a_{k}(0))(a_{k}-a_{k}(0))\Big{)}$
$=\sum_{k\in[n]}\Big{(}\frac{\phi^{\prime\prime}(a_{k}(0))}{2}+R(a_{k}-a_{k}(0))\Big{)}(a_{k}-a_{k}(0))^{2},$
where $R$ captures the error produced by Taylor approximation and satisfies $\lim_{z\rightarrow 0}R(z)=0$. Consider $L_{n}(\alpha)=\int\upsilon(w,b)\big{(}\alpha(w,b)\big{)}^{2}\mathrm{d}\mu_{n}$ where $\upsilon(w_{k}(0),b_{k}(0))=\frac{\phi^{\prime\prime}(a_{k}(0))}{2}+R(a_{k}-a_{k}(0))$ for $k\in[n]$. Hence, if $\bar{\boldsymbol{a}}$ solves (40), $\bar{\alpha}_{n}$, which satisfies $\bar{\alpha}_{n}(w_{k}(0),b_{k}(0))=\bar{a}_{k}-a_{k}(0)$ for $k\in[n]$, solves the following problem:
$\min_{\alpha\in\Theta_{\mathrm{ReLU}}}L_{n}(\alpha)\quad\ \mathrm{s.t.}\ \langle\alpha,\sigma\rangle_{n}=y_{i}-f(x_{i},\hat{\theta}).$
Let $\bar{\alpha}_{n}^{\prime}$ be the solution to the following problem:
${\min}_{\alpha\in\Theta_{\mathrm{ReLU}}}L_{n}^{\prime}(\alpha)\quad\mathrm{s.t.}\ \langle\alpha,\sigma_{i}\rangle_{n}=y_{i}-f(x_{i},\hat{\theta}),\ i\in[m],$
where $\text{where }L_{n}^{\prime}(\alpha)\triangleq\int\frac{\phi^{\prime\prime}(0)}{2}\big{(}\alpha(w,b)\big{)}^{2}\mathrm{d}\mu_{n}.$
Notice that
$|L_{n}(\alpha)-L_{n}^{\prime}(\alpha)|$ $=|\int\Big{(}\upsilon(w,b)-\frac{\phi^{\prime\prime}(0)}{2}\Big{)}\alpha^{2}(w,b)\mathrm{d}\mu_{n}|$
$\leq\max_{k\in[n]}\Big{|}\frac{\phi^{\prime\prime}(a_{k}(0))}{2}-\frac{\phi^{\prime\prime}(0)}{2}+R(a_{k}-a_{k}(0))\Big{|}\cdot\|\alpha\|_{n}^{2}.$
Following the same analysis as in the proof of Proposition 17, we have with high probability $\max_{k\in[n]}\bar{a}_{k}-a_{k}(0)=o_{p}(1)$, $\max_{k\in[n]}\bar{a}^{\prime}_{k}-a_{k}(0)=o_{p}(1)$, $\|\bar{\alpha}_{n}\|_{n}^{2}=O_{p}(1)$, and $\|\bar{\alpha}_{n}^{\prime}\|_{n}^{2}=O_{p}(1)$. Meanwhile, noticing that $\max_{k\in[n]}|a_{k}(0)|=O_{p}(\sqrt{(\log n)/n})$ holds with high probability and by the continuity of $\phi^{\prime\prime}$, we have $\max_{k\in[n]}|\phi^{\prime\prime}(a_{k}(0))-\phi^{\prime\prime}(0)|=o_{p}(1)$. Therefore, we have
$|L_{n}(\bar{\alpha}_{n})-L^{\prime}_{n}(\bar{\alpha}_{n})|,\ |L_{n}(\bar{\alpha}_{n}^{\prime})-L^{\prime}_{n}(\bar{\alpha}_{n}^{\prime})|=o_{p}(1).$
The remainder analysis proceeds as in the proof of Proposition 17.
So far, we have shown that the training converges to zero training error, the lazy training occurs, and the implicit bias in the infinite width limit can be exactly described as in Proposition 17. Then Proposition 22 directly applies and gives the claimed results.
@@ B.1 Linearization
$H(t)=\frac{1}{n}J_{\theta}\boldsymbol{f}(t)(\nabla^{2}\Phi(\theta))^{-1}J_{\theta}\boldsymbol{f}(t)^{T},$
@@
Consider a single-hidden-layer univariate network (1) with ReLU or absolute value activation function, $d$ input units and $n$ hidden units. Assume $n$ is sufficiently large. Consider parameter initialization (2). Assume $\mathcal{B}$ has a compact support. Given any finite training data set $\{(x_{i},y_{i})\}_{i=1}^{m}$ that satisfies $x_{i}\neq x_{j}$ when $i\neq j$ and $\{\|x_{i}\|_{2}\}_{i=1}^{m}\subset\operatorname{supp}(p_{\mathcal{B}})$, consider mirror flow (4) with learning rate $\eta=\eta_{0}/n$, $\eta_{0}>0$ and potential function $\Phi$. Assume $\Phi$ can be written in the form of $\Phi(\theta)=\frac{1}{n^{2}}\sum_{k=1}^{3n+1}\phi\big{(}n(\theta_{k}-\hat{\theta}_{k})\big{)}$, where $\phi$ is a real-valued $C^{3}$-smooth function on $\mathbb{R}$ and has strictly positive second derivatives everywhere. Assume there exists constant $C_{1}>0$ satisfying
$\frac{\kappa_{1}(C_{1})}{\kappa_{2}(C_{1})}\geq\frac{8\sqrt{m}(B_{x}+1)K\|\boldsymbol{f}(0)-\boldsymbol{y}\|_{2}}{\lambda_{0}}\cdot\frac{1}{C_{1}},$ (72)
where $\lambda_{0}=\operatorname{plim}_{n\rightarrow\infty}\lambda_{\min}(\frac{1}{n}J_{\theta}\boldsymbol{f}(0)J_{\theta}\boldsymbol{f}(0)^{T})$, $m$ is the size of the training data set, $B_{x}=\max_{i\in[m]}\|x_{i}\|_{2}$, and $K$ is a constant determined by $\mathcal{B}$ and $\mathcal{A}$. Then with high probability over random initialization,
@@
$\|\boldsymbol{f}(t)-\boldsymbol{y}\|_{2}\leq\exp\Big{(}-\frac{\kappa_{1}(C_{1})\lambda_{0}}{4}\eta_{0}t\Big{)}\|\boldsymbol{f}(0)-\boldsymbol{y}\|_{2},\ \forall t\geq 0.$
@@
[Proof of Proposition 27] To ease the notation, let $J(t)=J_{\theta}\boldsymbol{f}(t)$ and $Q(t)=(\nabla^{2}\Phi(\theta(t)))^{-1}$. For $i\in[m]$ and $k\in[n]$, let $\sigma_{ik}(t)=\sigma(w_{k}(t)^{T}x_{i}-b_{k}(t))$ and $\sigma^{\prime}_{ik}(t)=\sigma^{\prime}(w_{k}(t)^{T}x_{i}-b_{k}(t))$.
@@
$\begin{dcases}I_{1}=\frac{1}{n}\sum_{k=1}^{n}|\sigma_{ik}(t)\sigma_{jk}(t)-\sigma_{ik}(0)\sigma_{jk}(0)|;\\ I_{2}=|x_{i}^{T}x_{j}|\frac{1}{n}\sum_{k=1}^{n}|a_{k}(t)^{2}\sigma^{\prime}_{ik}(t)\sigma^{\prime}_{jk}(t)-a_{k}(0)^{2}\sigma^{\prime}_{ik}(0)\sigma^{\prime}_{jk}(0)|;\\ I_{3}=\frac{1}{n}\sum_{k=1}^{n}\Big{|}a_{k}(t)^{2}\sigma^{\prime}_{ik}(t)\sigma^{\prime}_{jk}(t)-a_{k}(0)^{2}\sigma^{\prime}_{ik}(0)\sigma^{\prime}_{jk}(0)|.\end{dcases}$ (74)
@@
$\frac{\mathrm{d}}{\mathrm{dt}}\|\boldsymbol{f}(t)-\boldsymbol{y}\|_{2}^{2}\leq-2\eta_{0}\frac{\kappa_{1}(C_{1})\lambda_{0}}{4}\|\boldsymbol{f}(t)-\boldsymbol{y}\|_{2}^{2}$
@@
$\|\boldsymbol{f}(t)-\boldsymbol{y}\|_{2}\leq\exp(-\frac{\kappa_{1}(C_{1})\lambda_{0}}{4}\eta_{0}t)\|\boldsymbol{f}(0)-\boldsymbol{y}\|_{2}.$ (77)
@@
$\|J_{a_{k}}\boldsymbol{f}(t)\|_{2}\leq\sqrt{m}\|J_{a_{k}}\boldsymbol{f}\|_{\infty}\leq\sqrt{m}(B_{x}\|w_{k}(t)\|_{2}+|b_{k}(t)|)\leq 2\sqrt{m}K(B_{x}+1).$ (78)
@@
$|\frac{\mathrm{d}}{\mathrm{dt}}a_{k}(t)|$ $\leq\frac{\eta_{0}}{n}\|Q(t)\|_{2}\|J_{a_{k}}\boldsymbol{f}(t)\|_{2}\|\boldsymbol{f}(t)-\boldsymbol{y}\|_{2}$ (79)
$\leq\frac{\eta_{0}}{n}\cdot\kappa_{2}(C_{1})\cdot 2\sqrt{m}K(B_{x}+1)\cdot\exp(-\frac{\kappa_{1}(C_{1})\lambda_{0}}{4}\eta_{0}t)\|\boldsymbol{f}(0)-\boldsymbol{y}\|_{2}.$
@@
$<\frac{8\kappa_{2}(C_{1})\sqrt{m}K(B_{x}+1)\|\boldsymbol{f}(0)-\boldsymbol{y}\|_{2}}{\kappa_{1}(C_{1})\lambda_{0}}n^{-1}$
@@
$\begin{dcases}\|J_{w_{k,r}}\boldsymbol{f}(t)\|_{2}\leq\sqrt{m}B_{x}|a_{k}(t)|\leq 2\sqrt{m}B_{x}K,\quad\forall k\in[n],r\in[d];\\ \|J_{b_{k}}\boldsymbol{f}(t)\|_{2}\leq\sqrt{m}|a_{k}(t)|\leq 2\sqrt{m}K,\quad\forall k\in[n];\\ \|J_{d}\boldsymbol{f}(t)\|_{2}=\sqrt{m}.\end{dcases}$ (81)
@@
$\|\tilde{\theta}(t)-\hat{\theta}\|_{\infty}\leq C_{1}n^{-1},\quad\|\tilde{\boldsymbol{f}}(t)-\boldsymbol{y}\|_{2}\leq\exp(-\frac{\kappa_{1}(C_{1})\lambda_{0}}{4}\eta_{0}t)\|\boldsymbol{f}(0)-\boldsymbol{y}\|_{2}.$
@@
$\kappa_{2}(C_{1})\kappa_{3}(C_{1})\Big{(}\lambda_{0}\frac{\kappa_{2}(C_{1})}{\kappa_{1}(C_{1})}+10m(B_{x}+1)^{2}K^{2}\Big{(}\frac{\kappa_{2}(C_{1})}{\kappa_{1}(C_{1})}\Big{)}^{2}\Big{)}\leq\frac{(\lambda_{0})^{2}}{16(B_{x}+1)K\sqrt{m}\|\boldsymbol{f}(0)-\boldsymbol{y}\|_{2}}.$ (82)
Then for any $x\in\mathbb{R}^{d}$ and with high probability over initialization,
@@
For any fixed $C>0$, consider $t_{1}=\inf\{t\geq 0\colon\iota(t)>C\}$. Now we aim to prove $t_{1}=+\infty$. For the sake of contradiction, assume $t_{1}<+\infty$. Then we have for $t\in[0,t_{1}]$, $\iota(t)\leq C$ and in particular, $\iota(t_{1})=C$.
@@
Next we bound the difference between $H(t)$ and $\tilde{H}(t)$. To simplify the notation, we let $\sigma_{ik}(t)=\sigma(w_{k}(t)^{T}x_{i}-b_{k}(t))$ for $i\in[m]$ and $k\in[n]$. For $i,j\in[m]$ we have the following decomposition:
$|H_{ij}(t)-\tilde{H}_{ij}(t)|\leq I_{1}+I_{2}+I_{3},$ (83)
@@
$\begin{dcases}I_{1}=\frac{1}{n}\sum_{k=1}^{n}\Big{|}\frac{\sigma_{ik}(t)\sigma_{jk}(t)}{\phi^{\prime\prime}(n(a_{k}(t)-a_{k}(0)))}-\frac{\sigma_{ik}(0)\sigma_{jk}(0)}{\phi^{\prime\prime}(n(\tilde{a}_{k}(t)-a_{k}(0)))}\Big{|};\\ I_{2}=\frac{1}{n}\sum_{k=1}^{n}\sum_{r=1}^{d}\Big{(}|x_{i,r}x_{j,r}|\Big{|}\frac{a_{k}(t)^{2}\sigma^{\prime}_{ik}(t)\sigma^{\prime}_{jk}(t)}{\phi^{\prime\prime}(n(w_{k,r}(t)-w_{k,r}(0)))}\Big{|}\Big{)}+\Big{|}\frac{a_{k}(t)^{2}\sigma^{\prime}_{ik}(t)\sigma^{\prime}_{jk}(t)}{\phi^{\prime\prime}(n(b_{k}(t)-b_{k}(0)))}\Big{|}\\ I_{3}=\frac{1}{n}\Big{|}\frac{1}{\phi^{\prime\prime}(n(d(t)-d(0)))}\Big{|}.\end{dcases}$ (84)